### v1

# Challenges for first-principles nuclear structure: 11Li and 29F

Abstract
Ab initio calculations of atomic nuclei have had many successes in recent years. Nonetheless, important challenges that resist even brute-force calculation remain. As archetypal examples of these challenges, we consider 11Li and 29F, well known halo nuclides situated on islands of inversion. The deformed intruder levels, which are primarily two-particle, two-hole neutron excitations with respect to naive spherical shell model configurations, are slow, with respect to increases of the model space, to take their rightful place as the lowest levels; in fact the spherical and deformed configurations likely mix strongly. We suggest these systems prototype the challenges for other important intruder states, and can serve as useful testbeds for potential approaches.
Introduction.
Simple models can be challenged by messy realities; adapting to reality can lead to new, more insightful models. As an early example of this, the spherical independent-particle shell-model of atomic nuclei was confronted by the reality of collective quadrupolar dynamics, which lead, first to the deformed Nilsson model [1, 2], and then to Elliott’s SU(3) picture [3, 4, 5, 6, 7, 8]. Particularly instructive is the history of the configuration-interaction (CI) shell-model, where one expands a many-body state in a basis, allowing the Schrödinger equation to be written as a matrix eigenvalue problem [9].
For practical considerations, one must truncate the many-body space. Starting from a mean-field picture, the spherical shell model assumes an energetically ordered set of single-particle orbitals, which determines the priority of filling; this is called the aufbau (building up) principle in atomic physics. Such an expected priority of filling naturally leads to the concept of a valence or active space, bounded from below by a frozen core of filled orbitals, and from above by empty, excluded orbitals. Valence space shell model calculations have been very successful in reproducing many features of low-lying excitation spectra as well as many electromagnetic and weak decays, and are still widely used today [10, 11].
Because of the strength of the nuclear force, one must take account of configurations that produce holes in the core, and those that catapult particles into the excluded orbitals. Originally this was done through perturbation theory [9, 12]. But careful analyses showed that such perturbative calculations would generically run into difficulties [13, 14].
To explain the difficulties, one must introduce the enveloping concept of intruders. In the simplest definition of intruders, one divides the infinite Hilbert space into the model space and the excluded space, represented by projection operators $P$ and $Q$, respectively, such that $P+Q=1$. Ideally, low-lying wave functions live mostly in the model space $P$; wave functions in the excluded space $Q$ are expected to lie higher in energy. An intruder upends this expectation: it is a wave function lying low in the spectrum, but dominated by components from the excluded space.
The existence of intruder states in nuclei was recognized early on; for example the realization [15] that the ground state of 11Be has $J^{\pi}=1/2^{+}$, whereas the naive shell model would suggest instead $1/2^{-}$. Some of best known intruders may be the primarily four-particle, four-hole (alpha cluster excitation) Hoyle state in 12[16, 17] and the related first excited $0^{+}$ state at 6.05 MeV in 16[18, 19], and $0^{+}$ states at 3.35 and 5.21 MeV in 40Ca [20]. The notorious islands of inversion [21, 22], where the ground state is itself an intruder relative to the valence space, as in 11Be  [15], provide yet more examples. Intruder states are often associated with shape coexistence, where spherical and deformed states lie close in energy and mix. The coexisting deformed states can be interpreted in terms of particle-hole excitations across shell boundaries and are found in many regions of the chart of the nuclides; see [23], especially their Figs. 2 and 8.
The term “intruder” implies something unnatural or unusual about a state, especially when contrasted against so-called “normal” states as we do below. Yet it can be argued that while these deformed particle-hole states are intruders from the viewpoint of spherical orbitals, they are in fact are naturally expected by applying the aufbau principle to a basis of deformed single-particle orbitals such as the Nilsson picture [24, 25, 26, 27]. Indeed, soon after it was suggested that 11Li has a neutron halo, it was pointed out this could be easily explained in the Nilsson model [24].
But intruders or, at least, states which are intruders in a spherical shell model picture, cause havoc in perturbative approaches, which assume states from the excluded space are far away in energy. The failure of perturbation theory in the face of intruders caused an epistemological crisis [28] in nuclear shell model approaches: if perturbative expansions failed, how could one rigorously derive shell-model interactions from microscopic data such as nucleon-nucleon scattering phase shifts, rather than fitting to many-body spectra? Because of this, the nuclear shell model limped on primarily as an empirical/phenomenological method for nearly two decades. (A modern alternative to perturbation theory is the in-medium similarity renormalization group (IMSRG) [29]. Despite being nonperturbative, IMSRG also faces complications when dealing with intruder states [30].)
It was the introduction of the no-core shell model [9, 31] (NCSM) that revived nuclear configuration-interaction as a potent tool with a rigorous ab initio basis. Because the NCSM from the very beginning includes, in principle, all excitations, and those excitations are handled nonperturbatively, the original intruder crisis went away (or, to put it another way, since there is no core, there can be no particle-hole excitations to cause trouble).
No-core shell model basis states themselves are constructed from harmonic oscillator single-particle states described by the oscillator frequency $\omega$, or more commonly, the oscillator energy parameter $\hbar\omega$. A key label of many-body configurations is the number $N_{\text{ex}}$} of harmonic oscillator excitation quanta relative to the lowest filling of oscillator shells. Configurations of a given $N_{\text{ex}}$} are sometimes termed $N_{\mathrm{ex}}\hbar\omega$ configurations: the lowest configurations, with $N_{\mathrm{ex}}=0$, are $0\hbar\omega$ configurations, etc. A given model space is then labeled by $N_{\mathrm{max}}$, the maximum excitation allowed, or, alternately, as an $N_{\mathrm{max}}\hbar\omega$ space. Thus, while there is formally no valence space in the NCSM, it is nonetheless convenient to refer to the $N_{\mathrm{max}}=0$ or $0\hbar\omega$ space as the naive valence space. Then we define states dominated by configurations in the naive valence space as “normal” states, and states dominated by configurations outside the naive valence space, in this scheme, as “intruders.” Applying the concept of intruders to the NCSM is not new [32, 33, 34, 35, *caprio2022:10be-shape-sdanca21, 37, 38, 39].
Figure 1: Schematic of normal (left side) versus two-particle, two-hole intruder (right side) configurations. Dashed lines represent spherical shell closures. For our example of 11Li, the filled proton orbitals (diagonal fill) is $0s$ while the filled neutron orbitals (solid fill) consist of 0s0s-$0p$. For 29F, the filled proton and neutron orbitals are 0s0s-$0p$ and 0s0s-0p0p-$1s0d$, respectively.
In this paper, we discuss detailed NCSM calculations of two cases from the islands of inversion, 11Li and 29[22, 40]. These two exemplars are parallel, in that in the naive spherical shell model, one would expect the lowest configuration to consist of filled neutron oscillator shells ($0s$ and $0p$ in the case of 11Li, and $0s$, $0p$, and $1s0d$ for 29F) and a single proton outside a filled shell. Experimental evidence, however, supports the ground state having a significant contribution from two-particle, two-hole neutron excitations. Fig. 1 provides a schematic illustration. Our calculations highlight significant differences between the normal and intruder states: the normal states are nearly spherical, while the intruder states are strongly deformed.
Furthermore, while the ground states are known to be the intruder (or, more accurately, a strong mixture of the normal and intruder state), in the NCSM calculations, the intruder starts high in energy relative to the lowest normal state, for small model spaces, and only slowly, with respect to the model space parameter $N_{\mathrm{max}}$, comes down in energy. Thus, rather than the failure of perturbation theory as in the earlier intruder crisis, the problem with intruder states in the NCSM is their slow convergence with the model space size, relative to normal states. This relative misplacement of the ‘true’ ground state, and the slow convergence, provides a challenge to current first principles nuclear structure theory.
No-core shell model calculations.
Our NCSM calculations [9, 31] build upon the occupation representation/second quantization, expanding wave functions in a basis of Slater determinants with fixed total $z$-component of angular momentum $J_{z}$ or $M$ (hence, ‘M-scheme basis’). The official practice of NCSM calculations is to go to large enough $N_{\mathrm{max}}$ so as to converge to the $N_{\mathrm{max}}\rightarrow\infty$ limit, or nearly so; in large part this paper examines the consequences of not being able to do so.
We use two sets of interaction matrix elements. The first is derived from the Entem-Machleidt next-to-next-to-next-to-leading-order (N3LO) chiral effective theory [41], softened by similarity-group renormalization [42, 43] (SRG) evolution to a resolution parameter value of $\lambda=2.0\,{\mathrm{fm}}^{-1}$. We do not include three-body forces. The second interaction, Daejeon16 [44], is phase-shift-equivalent to the Entem-Machleidt interaction, but adjusted to reproduce the binding energies (and some excitation energies) in several nuclei with $A\leq 16$, without a three-body force. The latter is known to be “softer” than the former, since it has been more drastically SRG evolved, to $\lambda=1.5\,{\mathrm{fm}}^{-1}$. This leads to faster convergence of energies with increase of the model space, as we see in our calculations. For our observables such as the quadrupole moment and radius, we use the bare (unrenormalized) operators and charges. We choose the value of $\hbar\omega$ for each nuclide and interaction by approximately minimizing the lowest energy (ostensibly the ground state energy, although in our calculations the lowest state is predominantly “normal”, while the true ground state would in fact be predominantly “intruder”). NCSM calculations were carried out using the codes BIGSTICK [45] and MFDn [46, *shao2018:ncci-preconditioned].
To gain insight into the structure of the wave functions, we decompose via the Lanczos algorithm [48, 49, 50, 36], into irreps of Elliott’s SU(3) [51, 36, 39]. These irreps can be in turn approximately interpreted in terms of deformation parameters $\beta$ and $\gamma$ [52]. We can also simultaneously decompose into $N_{\mathrm{ex}}$, the number of oscillator excitations.
Refer to captionRefer to caption
Figure 2: (Left) Relative energies for the first two $1/2^{-}$ and $3/2^{-}$ levels, and first $5/2^{-}$ and $7/2^{-}$ levels, of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, as a function of $N_{\mathrm{max}}$, for (a) a chiral N3LO interaction and (b) the Daejeon16 interaction. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed basis parameter of ${\hbar\omega}=18\,{\mathrm{MeV}}$ for N3LO and ${\hbar\omega}=12.5\,{\mathrm{MeV}}$ for Daejeon16. The shading of the symbols indicates the structure, normal (shaded) or intruder (open), which might naively be expected for the level given the energy evolution, although the first two $1/2^{-}$ levels are undergoing an avoided crossing at higher $N_{\text{max}}$} in the Daejeon16 calculations (see text). (Right) Decompositions by $N_{\text{ex}}$} for first two $3/2^{-}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, in a calculation with ${N_{\text{max}}}=8$, for the Daejeon16 interaction (with ${\hbar\omega}=12.5\,{\mathrm{MeV}}$).
Case study: $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}.
11Li was an early and dramatic exemplar of halo nuclei [53]. Neutron knockout experiments show nearly equal contributions of neutron $1s_{1/2}$ and $0p_{1/2}$ orbitals [54], providing strong evidence for a ground state comprised of a nearly equal mix of normal and intruder states. 11Li is well known to be a challenge to describe fully microscopically, i.e., in terms of three protons and eight neutrons, for example in the NCSM [55, 56, 57], or using fermionic molecular dynamics (FMD) [58]. The FMD calculation addresses mixing of $(p_{1/2})^{2}$ and $(s_{1/2})^{2}$ configurations, that is, normal and intruder states. In many other calculations, 11Li has been approximated as a three-body system, typically $^{9}\mathrm{Li}+2n$ [59, 60, 61, 62, 63, 64, 65, 66, 67, 68], also allowing $(p_{1/2})^{2}$-$(s_{1/2})^{2}$ mixing to be investigated.
Our calculated excitation spectrum is shown in Fig. 2, as a function of $N_{\mathrm{max}}$ for both the chiral N3LO and Daejeon16 interactions (at fixed basis parameter ${\hbar\omega}=18\,{\mathrm{MeV}}$ and $12.5\,{\mathrm{MeV}}$, respectively). In the rightmost panels, we display the $N_{\text{ex}}$} decomposition of the ground state and excited $3/2^{-}$ states, in particular, illustrating how we may identify normal and intruder states from their NCSM wave functions [35, 37, 39].
The $0{\hbar\omega}$ shell model picture for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} gives just $3/2^{-}$ and $1/2^{-}$ states, corresponding to finding the valence neutron in the $p_{3/2}$ and $p_{1/2}$ orbitals, respectively. Indeed, the two lowest calculated states (shaded symbols in Fig. 2) match this expectation, and, from their decomposition into configurations with different $N_{\text{ex}}$}, they are seen to be normal (with an exception related to mixing at high $N_{\text{max}}$}, as discussed below). This is illustrated for the $3/2^{-}$ ground state in Fig. 2 (bottom right). Namely, the single largest contribution to the probability comes from ${N_{\text{ex}}}=0$ (in this case, a single configuration, the filled neutron $0p$ shell), although, naturally for a no-core calculation (see, e.g., Fig. 8 of Ref. [37]) this $0{\hbar\omega}$ configuration is “dressed” with a tail of smaller contributions from excited configurations.
Then, several “intruder” states arise at higher energy in the calculation (open symbols in Fig. 2), starting with a closely spaced $3/2^{-}$ and $1/2^{-}$. (There are many more such states, of which we only show the first $7/2^{-}$ and $5/2^{-}$ as representative examples.) While these start at high energy, for low $N_{\text{max}}$}, they rapidly descend with increasing $N_{\text{max}}$}. The intruder nature of the excited $3/2^{-}$ and $1/2^{-}$ states (again, with an exception related to mixing, as discussed below) is illustrated for the excited $3/2^{-}$ state in Fig. 2 (top right). Note the negligible $0{\hbar\omega}$ contribution, of a few percent, while the strongest contribution comes from $2{\hbar\omega}$ configurations, which are again dressed with a tail of more highly excited oscillator configurations.
Because Daejeon16 is a “softer” interaction, it is perhaps unsurprising that the intruder states found in the Daejeon16 calculation [Fig. 2(b)] descend in energy much sooner (with increasing $N_{\mathrm{max}}$) than those found in the N3LO calculation [Fig. 2(a)]. They also start to level off in energy, suggesting they are approaching the values they would converge to in an untruncated calculation.
As the energies of the intruder states approach those of the normal states, of the same angular momentum and parity, these states can mix. In the Daejeon16 calculations [Fig. 2(b)], the normal and intruder $1/2^{-}$ states (upward triangles) approach in energy first (due to the high excitation energy of the normal $1/2$- state, relative to the spacing between the intruder $3/2^{-}$ and $1/2^{-}$ states). They approach to within $\sim 2\,{\mathrm{MeV}}$, with increasing $N_{\text{max}}$}, but then get no closer. The excited $1/2^{-}$ state (open upward triangle) initially tracks the other intruder states, at lower $N_{\text{max}}$}, but then appears to hit a floor, at lower $N_{\text{max}}$}, while the other intruder states continue to fall in energy. On the other hand, the first $1/2^{-}$ state (shaded upward triangle), while stable in excitation energy below ${N_{\text{max}}}=4$, is pushed down in excitation energy, for higher $N_{\text{max}}$}. This behavior is suggestive of the level repulsion arising in an avoided crossing (e.g., Ref. [69]). Indeed, the $N_{\text{ex}}$} decompositions (not shown) of the wave functions for the calculated $1/2^{-}$ states suggest that the “normal” and “intruder” states are nearly maximally mixed by ${N_{\text{max}}}=12$. (The mixing matrix element between the normal and intruder states can thus be deduced to be on the order of half the minimal separation, or $\sim 1\,{\mathrm{MeV}}$.) With further increasing $N_{\text{max}}$}, we may expect that the intruder $1/2^{-}$ would emerge from the other side of the avoided crossing and resume tracking the energies of the other intruder states, thus supplanting the normal $1/2^{-}$ state as the lowest $1/2^{-}$ state.
Although the normal and intruder $3/2^{-}$ states [squares in Fig. 2(b)] do not yet approach as closely at the $1/2^{-}$ states in these calculations, the $N_{\text{ex}}$} decompositions of the wave functions (not shown) suggest about $10\%$ mixing already by ${N_{\text{max}}}=12$. We can anticipate that the mixing will continue to increase, as the intruder state continues to approach the ground state. How strongly the states will mix, and whether or not they might actually cross, depends in detail on the energy convergence and the strength of the mixing matrix element for the $3/2^{-}$ states, so we do not attempt a quantitative analysis here. We simply wish to highlight the importance of the intruder state and mixing in understanding even the ground state of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} in an ab initio calculation.
Figure 3: Dimensionless ratio $Q/r^{2}$ for first $3/2^{-}$ level (shaded symbols) and second $3/2^{-}$ level (open symbols) of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, for the protons (left) and neutrons (right), calculated with the chiral N3LO (top) and Daejeon16 (bottom) interactions. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed $\hbar\omega$} (as indicated). The experimental value for $Q_{p}/r_{p}^{2}$ [70, 71] for the $3/2^{-}$ ground state is shown for comparison (solid square).
Refer to captionRefer to caption
Figure 4: Decompositions by $\mathrm{SU}(3)$, jointly with the number of oscillator quanta, for the first two $3/2^{-}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, calculated with the chiral N3LO (top) and Daejeon16 (bottom) interactions, for $N_{\mathrm{max}}=8$ and ${\hbar\omega}=12.5\,{\mathrm{MeV}}$. These are shown as histograms (top) and arranged by the Bohr deformation variables corresponding to the given $\mathrm{SU}(3)$ quantum numbers (bottom). Contributions from the $0{\hbar\omega}$ (shaded circles) and $2{\hbar\omega}$ (open circles) spaces are shown. Contributions from irreps which are degenerate with respect to the $\mathrm{SU}(3)$ Casimir operator (connected by dotted lines) cannot be distinguished (and, for plotting purposes, the contribution has, arbitrarily, been distributed equally between these irreps).
Now we turn to the question of the role of deformation in the intruder states, and thus (after mixing) in the $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} halo ground state. While deformation itself is not directly observable, the electric quadrupole moment, an observable, is strongly sensitive to deformation. In Fig. 3, we show separately the proton and neutron quadrupole moments, or, more precisely, dimensionless quantities obtained by scaling the quadrupole moments to the mean-square proton and neutron radii, respectively. Such a dimensionless ratio $Q/r^{2}$ provides a measure of the quadrupole deformation, and (at least, for axially symmetric rotational nuclei) may be approximately related to the Bohr deformation variable $\beta$ [see, e.g., Eq. (13) of Ref. [72]]. From a practical viewpoint, it is also more rapidly convergent in the NCSM calculations than either the quadrupole moment or mean square radius individually [57, 72].
Take first the proton ratio $Q_{p}/r_{p}^{2}$ [Fig. 3 (left)]. Experimental measurements of the quadrupole moment are, of course, only sensitive to this proton contribution. (The experimental quadrupole moment for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} shows only a slight increase in magnitude relative to $^{\makebox[3.98613pt][r]{$\scriptstyle 9$}}_{\makebox[3.98613pt][r]{$\scriptstyle$}}\mathrm{Li}$} [73, 70].) Note that the calculated quadrupole moment is negative, for both the ground and excited $3/2^{-}$ states, so increasing magnitude goes downward in the plot. For the ground state, a negative sign for the quadrupole moment is as expected for a single nucleon outside a closed shell (e.g., Refs. [74, 75]), while, for the excited $3/2^{-}$, one might associate a negative quadrupole moment with prolate deformation in a rotational $K=1/2$ band (e.g., Fig. 1 of Ref. [76]). Regardless, the proton $Q/r^{2}$ is nearly the same for the normal and intruder states; this at least suggests the proton structure is not very different between the two. Our results are in close agreement with experiment [70, 71], shown as a solid square in Fig. 3, in the calculations with either the N3LO or Daejeon16 interaction. (Previous NCSM calculations obtained similar values of $Q_{p}/r_{p}^{2}$ [55, 56], including prior calculations with the Daejeon16 interaction [57].)
Now consider the calculated neutron quadrupole moment or, rather, $Q_{n}/r_{n}^{2}$, [Fig. 3 (right)]. It is very small for the normal state ($Q_{n}/r_{n}^{2}\approx-0.2$), consistent with small deviations from a spherical filled neutron shell, but much larger for the intruder state ($Q_{n}/r_{n}^{2}\approx-1.3$), as one might expect if this is a highly deformed state arising from $2{\hbar\omega}$ excitations of neutrons across the $N=8$ shell closure. If the intruder $3/2^{-}$ state is taken to be a member of a $K=1/2$ band (with inverted spin orderings due to Coriolis staggering), a plausible speculation but likely an oversimplification given the likely role of triaxiality (discussed below), then the calculated ratio corresponds to a hefty prolate intrinsic deformation of $\beta_{n}\approx 0.6$ for the neutrons. At least, this is the behavior of the observables for the N3LO calculations [Fig. 3 (b)].
However, in the Daejeon16 calculations [Fig. 3 (d)], the situation is more subtle, as we might expect in the presence of mixing. For the ground state, the neutron quadrupole moment starts small, at low $N_{\text{max}}$}, as in the N3LO calculation, but then it starts steadily climbing, from ${N_{\text{max}}}=4$ onward. For the excited $3/2^{-}$ state, the neutron quadrupole moment attains a large value at low $N_{\text{max}}$}, as in the N3LO calculation, but then turns over and starts declining in magnitude by ${N_{\text{max}}}=10$. This reflects an intruder component (with large neutron deformation) mixing into the ground state, and a normal component (with negligible neutron deformation) mixing into the excited state. As the intruder $3/2^{-}$ continues to fall in energy, and mixing with the ground state becomes more pronounced, we can only expect this trend to continue, so that the $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} ground state structure arises from mixing with an intruder, introducing a deformed neutron halo.
To proceed beyond the simple measure of deformation provided by $Q/r^{2}$ and more properly characterize the “shape” of the state, we decompose the wave function into components in irreps of Elliott’s SU(3) [51] and interpret them approximately in terms of the standard quadrupole shape parameters $\beta$ and $\gamma$ [52]. In Fig. 4 we display the structure of the normal and intruder $3/2^{-}$ wave functions, first in terms of excitation quanta $N_{\mathrm{ex}}$, and then, within each ${N_{\text{ex}}}\hbar\omega$ space, by SU(3) irreps, and thus location on the $(\beta,\gamma)$ plane. (Specifically, these results are for the $N_{\mathrm{max}}=8$ calculation using Daejeon16 at $\hbar\omega=12.5$ MeV; the plots for higher values of $N_{\mathrm{max}}$ are not very different.) As already noted (Fig. 2), the “normal” $3/2^{-}_{1}$ state (in this calculation) is largely $0{\hbar\omega}$, while the $0{\hbar\omega}$ contribution to the “intruder” $3/2^{-}_{2}$ is less than $5\%$.
Moreover, the shapes of these states differ significantly. The normal $3/2^{-}_{1}$ state is dominated by a nearly spherical $N_{\mathrm{ex}}=0$ (1,0) irrep of SU(3) (filled circle in the $\beta$-$\gamma$ plot, where the area of the circle is proportional to the probability), while the intruder $3/2_{2}^{-}$ state has its largest contributions from highly deformed ${N_{\text{ex}}}=2$ $(5,2)$ and $(6,0)$ irreps (open circles in $\beta$-$\gamma$ plot).
Thus, despite the numerically large calculation and a priori complex nature of the microscopic wave functions, the final picture is fairly simple: the normal state follows with the naive spherical shell-model expectations, while the intruder is both mostly two-particle, two-hole and well-deformed.
Refer to captionRefer to caption
Figure 5: (Left) Relative energies for the lowest $1/2+$ and $5/2+$ levels (shaded symbols) and lowest “intruder” $1/2+$, $5/2+$, and $5/2+$ levels (open symbols), of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, as a function of $N_{\mathrm{max}}$, for (a) a chiral N3LO interaction and (b) the Daejeon16 interaction. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed $\hbar\omega$} (as indicated). (Right) Decompositions by $N_{\text{ex}}$} for first two $5/2^{+}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, in a calculation with ${N_{\text{max}}}=6$, for the Daejeon16 interaction (with ${\hbar\omega}=17.5\,{\mathrm{MeV}}$).
Figure 6: Dimensionless ratio $Q/r^{2}$ for the first $5/2^{+}$ level (shaded symbols) and second $5/2^{+}$ level (open symbols) of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, for the protons (left) and neutrons (right), calculated with the Daejeon16 interaction. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed $\hbar\omega$} (as indicated).
Refer to captionRefer to caption
Figure 7: Decompositions by $\mathrm{SU}(3)$, jointly with the number of oscillator quanta, for the first two $5/2^{+}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, calculated with the Daejeon16 interaction, for $N_{\mathrm{max}}=4$ and ${\hbar\omega}=17.5\,{\mathrm{MeV}}$. These are shown as histograms (top) and arranged by the Bohr deformation variables corresponding to the given $\mathrm{SU}(3)$ quantum numbers (bottom). Contributions from the $0{\hbar\omega}$ (shaded circles) and $2{\hbar\omega}$ (open circles) spaces are shown. Contributions from irreps which are degenerate with respect to the $\mathrm{SU}(3)$ Casimir operator (connected by dotted lines) cannot be distinguished (and, for plotting purposes, the contribution has, arbitrarily, been distributed equally between these irreps).
Case study: $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}.
29F may be viewed as a cousin to 11Li, shifted up by a major oscillator shell. It is another halo nucleus [77] on an island of inversion [78]. Based at least on model arguments, though not direct experimental evidence, it is also thought to be deformed [79, 80]. Our maximum model space is only $N_{\mathrm{max}}=6$, with an $M$-scheme dimension of $4.7\times 10^{9}$. Nonetheless we can see trends parallel to those in 11Li.
The energy levels and the $N_{\mathrm{ex}}$ distributions are shown in Fig. 5, with the chiral N3LO interaction calculated at $\hbar\omega=18$ MeV and Daejeon16 interaction at $\hbar\omega=17.5$ MeV. Similar to 11Li, the intruder states calculated with the softer Daejeon16 interaction are lower in energy relative to the normal states than those calculated with the N3LO interaction, and the normal and intruder excited $1/2^{+}$ states (upward triangles) already approach to within $\approx 2\,{\mathrm{MeV}}$. However, we do not yet see as strong a mixing between normal and intruder states, for the model spaces we can reach in these calculations.
The $J^{\pi}=5/2^{+}$ proton and neutron quadrupole moments, scaled by the respective mean-squared radii, are shown in Fig. 6, with open symbols for the intruder states and filled symbols for normal states. We show results here only using the Daejeon16 wave functions. The pattern is, again, very similar to that of 11Li: very little difference between the normal and intruder proton quadrupole moments, suggesting the proton deformation is nearly the same, while the normal neutron quadrupole moment is nearly zero, as expected for a filled spherical shell, and large for the intruder state, suggesting strongly deformed.
The picture of a spherical normal state and a deformed intruder is backed up in Fig. 7 by the decomposition into SU(3) irreps and interpretation via the Bohr deformation parameters. The intruder decomposition and deformation is more fragmented than for 11Li; this is unsurprising considering the lore that spin-orbit splitting, which breaks SU(3), increases with mass number $A$.
The main difference from 11Li is that taking 29F to higher $N_{\mathrm{max}}$, closer to convergence, is beyond current capabilities. It may not be necessary: we have strong evidence of both the importance and the numerical challenge of intruders for these island-of-inversion nuclides. Instead, using what we have learned, it may be time to turn to alternate many-body methods. Furthermore, we suggest these nuclides as excellent tests for new calculations.
Summary.
We have carried out large no-core shell model calculations of the island-of-inversion nuclides 11Li and 29F, using two interactions, a chiral N3LO interation and the phase-shift-equivalent interaction Daejeon16. In a naive spherical shell model picture, these nuclides should have very simple structure: spherical filled neutron shells and one proton outside filled proton shells. Instead, the ground state and other low-lying states are inverted to be primarily, or at least strongly mixed with, deformed two-particle, two-hole intruders. While the intruders descend towards the spherical normal states with increasing model space ($N_{\mathrm{max}}$), convergence is slow, even for the softer Daejeon16 interaction, and thus are a challenge to reproduce in the NCSM. Interestingly, the intruder states may be “simple” in a group-theoretical framework, that is, dominated by a few irreps.
Confronted with the paradox of the simple naive spherical picture versus the more complicated deformed reality, at least for the spherical shell model, we propose these nuclides as important test cases for advanced nuclear structure methods. Examples include the symmetry-adapted configuration-interaction calculations [81, 82, 83, 84], the recently introduced angular-momentum-projected coupled cluster method [85, 86], or perhaps novel methods, yet to be proposed, in which these “intruder” configurations become the new normal.
Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Numbers DE-FG02-95ER40934 and DE-FG02-03ER41272. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231, using NERSC awards NP-ERCAP0023497 and NP-ERCAP0031993. Part of this research was enabled by computational resources supported by a generous gift to SDSU from John Oldham.