### v2

Abstract
Ab initio calculations of atomic nuclei have had many successes in recent years. Nonetheless, important challenges that resist even brute-force calculation remain. As archetypal examples of these challenges, we consider 11Li and 29F, well known halo nuclides situated on islands of inversion. The deformed intruder levels, which are primarily two-particle, two-hole neutron excitations with respect to naive spherical shell model configurations, are slow, with respect to increases of the model space, to take their rightful place among, and potentially mix with, the lowest levels. We suggest these systems prototype the challenges for other important intruder states, and can serve as useful testbeds for potential approaches.
Introduction.
Simple models can be challenged by messy realities; adapting to reality can lead to new, more insightful models. As an early example of this, the spherical independent-particle shell-model of nuclei was confronted by collective quadrupolar dynamics, which lead, first to the deformed Nilsson model [1, 2], and then to Elliott’s $\mathrm{SU}(3)$ picture [3, 4, 5, 6, 7, 8]. Particularly instructive is the history of the configuration-interaction shell-model, where one expands a many-body state in a basis, allowing the Schrödinger equation to be written as a matrix eigenvalue problem [9].
For practical considerations, one must truncate the many-body space. Starting from a mean-field picture, the spherical shell model assumes an energetically ordered set of single-particle orbitals, which determines the priority of filling; this is called the aufbau (building up) principle in atomic physics. Such an expected priority of filling naturally leads to the concept of a valence or active space, bounded from below by a frozen core of filled orbitals, and from above by empty, excluded orbitals. Valence space shell model calculations have been very successful in reproducing many features of low-lying excitation spectra as well as many electromagnetic and weak decays, and are still widely used today [10, 11].
Because of the strength of the nuclear force, one must take account of those configurations with holes in the core, and those with particles catapulted into the excluded orbitals. Originally this was done through perturbation theory [9, 12]. But careful analyses showed that such perturbative calculations would generically run into difficulties [13, 14].
To explain the difficulties, one must introduce the concept of intruders. In the simplest definition of intruders, one divides the infinite Hilbert space into the model space and the excluded space, represented by projection operators $P$ and $Q$, respectively, such that $P+Q=1$. Ideally, low-lying wave functions live mostly in the model space $P$; wave functions in the excluded space $Q$ are expected to lie higher in energy. An intruder upends this expectation: a state lying low in the spectrum, but dominated by components from the excluded space.
The existence of intruder states in nuclei was recognized early on: for example the realization [15] that the ground state of 11Be has $J^{\pi}=1/2^{+}$, whereas the naive shell model would suggest $1/2^{-}$. Some of the best known intruders are the four-particle, four-hole (alpha cluster excitation) Hoyle state in 12[16, 17] and the related first excited $0^{+}$ state at 6.05 MeV in 16[18, 19], and $0^{+}$ states at 3.35 and 5.21 MeV in 40Ca [20]. The notorious islands of inversion [21, 22], where the ground state is itself an intruder relative to the valence space, as in 11Be [15], provide more examples. Intruder states are often associated with shape coexistence, where spherical and deformed states lie close in energy. The coexisting deformed states can be interpreted in terms of particle-hole excitations across shell boundaries and are found in many regions of the chart of the nuclides (see Ref. [23], especially their Figs. 2 and 8).
The term “intruder” implies something unnatural or unusual about a state, especially when contrasted against so-called “normal” states as we do below. Yet while these deformed particle-hole states are intruders from the viewpoint of spherical orbitals, they are in fact are naturally expected by applying the aufbau principle to a basis of deformed single-particle orbitals such as the Nilsson picture [24, 25, 26, 27]. Indeed, soon after it was suggested that 11Li has a neutron halo, it was pointed out this could be easily explained in the Nilsson model [24].
But intruders or, at least, states which are intruders in a spherical shell model picture, cause havoc in perturbative approaches, which assume states from the excluded space are far away in energy. The failure of perturbation theory in the face of intruders caused an epistemological crisis [28] in nuclear shell model approaches: if perturbative expansions failed, how could one rigorously derive shell-model interactions from microscopic data such as nucleon-nucleon scattering phase shifts, rather than fitting to many-body spectra? Because of this, the nuclear shell model limped on primarily as an empirical/phenomenological method for nearly two decades. (A modern alternative to perturbation theory is the in-medium similarity renormalization group (IMSRG) [29]. Despite being nonperturbative, IMSRG also faces complications when dealing with intruder states [30].)
It was the introduction of the no-core shell model [9, 31] (NCSM) that revived nuclear configuration-interaction as a potent tool with a rigorous ab initio basis. Because the NCSM from the very beginning includes, in principle, all excitations, and those excitations are handled nonperturbatively, the original intruder crisis went away (or rather, since there is no core, there can be no particle-hole excitations to cause trouble).
No-core shell model basis states themselves are constructed from harmonic oscillator single-particle states described by the oscillator frequency $\omega$, or more commonly, the oscillator energy parameter $\hbar\omega$. A key label of many-body configurations is the number $N_{\text{ex}}$} of harmonic oscillator excitation quanta relative to the lowest filling of oscillator shells. Configurations of a given $N_{\text{ex}}$} are sometimes termed $N_{\mathrm{ex}}\hbar\omega$ configurations: the lowest configurations, with $N_{\mathrm{ex}}=0$, are $0\hbar\omega$ configurations, etc. A given model space is then labeled by $N_{\mathrm{max}}$, the maximum excitation allowed, or, alternately, as an $N_{\mathrm{max}}\hbar\omega$ space. Thus, while there is formally no valence space in the NCSM, it is nonetheless convenient to refer to the $N_{\mathrm{max}}=0$ or $0\hbar\omega$ space as the naive valence space. Then we define states dominated by configurations in the naive valence space as “normal” states, and states dominated by configurations outside the naive valence space, in this scheme, as “intruders.” Applying the concept of intruders to the NCSM is not new [32, 33, 34, 35, 36, 37, 38, 39].
Figure 1: Schematic of normal (left side) versus two-particle, two-hole intruder (right side) configurations. Dashed lines represent spherical shell closures. For our example of 11Li, the filled proton orbital (hashed fill) is $0s$ while the filled neutron orbitals (solid fill) consist of 0s0s-$0p$. For 29F, the filled proton and neutron orbitals are 0s0s-$0p$ and 0s0s-0p0p-$1s0d$, respectively.
In this paper, we discuss detailed NCSM calculations of two cases from the islands of inversion, 11Li and 29[22, 40, 41]. These two exemplars are parallel, in that in the naive spherical shell model, one would expect the lowest configuration to consist of filled neutron oscillator shells ($0s$ and $0p$ in the case of 11Li, and $0s$, $0p$, and $1s0d$ for 29F) and a single proton outside a filled shell. Experimental evidence, however, supports the ground state having a significant contribution from two-particle, two-hole neutron excitations. Fig. 1 provides a schematic illustration. Our calculations highlight significant differences between the normal and intruder states: the normal states are nearly spherical, while the intruder states are strongly deformed.
Furthermore, while the ground states are known to be the intruder (or, more accurately, a strong mixture of the normal and intruder state), in the NCSM calculations, the intruder starts high in energy relative to the lowest normal state, for small model spaces, and only slowly, with respect to the model space parameter $N_{\mathrm{max}}$, comes down in energy. Thus, rather than the failure of perturbation theory as in the earlier intruder crisis, the problem with intruder states in the NCSM is their slow convergence with the model space size, relative to normal states. This relative misplacement of the ‘true’ ground state, and the slow convergence, provides a challenge to current first principles nuclear structure theory.
No-core shell model calculations.
Our NCSM calculations [9, 31] build upon the occupation representation/second quantization, expanding wave functions in a basis of Slater determinants with fixed total $z$-component of angular momentum $J_{z}$ or $M$ (hence, ‘M-scheme basis’). The official practice of NCSM calculations is to go to large enough $N_{\mathrm{max}}$ so as to converge to the $N_{\mathrm{max}}\rightarrow\infty$ limit, or nearly so; in large part this paper examines the consequences of not being able to do so.
We use two sets of interaction matrix elements. The first is derived from the Entem-Machleidt next-to-next-to-next-to-leading-order (N3LO) chiral effective theory [42], softened by similarity-group renormalization [43, 44] (SRG) evolution to a resolution parameter value of $\lambda=2.0\,{\mathrm{fm}}^{-1}$. We do not include three-body forces. The second interaction, Daejeon16 [45], is again derived from the Entem-Machleidt interaction, but in this case more drastically softened by SRG evolution, to $\lambda=1.5\,{\mathrm{fm}}^{-1}$ (after which phase-shift-equivalent transformations are applied to reproduce the binding energies, and selected excitation energies, in several nuclei with $A\leq 16$, without including any induced three-body force). The resulting “softer” interaction leads to faster convergence of energies with increase of the model space, as we see in our calculations. For our observables, namely, the quadrupole moment and radius, we use the bare (unrenormalized) operators and charges.
We choose the value of $\hbar\omega$ for each nuclide and interaction by approximately minimizing the lowest energy (ostensibly the ground state energy, although in our calculations the lowest state is predominantly normal, while the true ground state would in fact be predominantly intruder). NCSM calculations were carried out using the codes BIGSTICK [46] and MFDn [47, 48].
To gain insight into the structure of the wave functions, we decompose them, via the Lanczos algorithm [49, 50, 51, 36], into irreps of Elliott’s $\mathrm{SU}(3)$ [51, 36, 39]. These irreps can be approximately interpreted in terms of deformation parameters $\beta$ and $\gamma$ [52]. We also simultaneously decompose into $N_{\mathrm{ex}}$, the number of oscillator excitations.
Figure 2: Relative energies for the normal $1/2^{-}$ and $3/2^{-}$ (blue, shaded symbols) and first intruder $1/2^{-}$ through $7/2^{-}$ (red, open symbols) levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, for (a) a chiral N3LO interaction and (b) the Daejeon16 interaction, shown as functions of $N_{\text{max}}$} (at fixed $\hbar\omega$}, as indicated). Although states are designated as normal or intruder in this figure according to what might naively be expected for the level given the energy evolution, the first two $1/2^{-}$ levels start to undergo an avoided crossing at higher $N_{\text{max}}$} in the Daejeon16 calculations (see text).
Refer to captionRefer to caption
Figure 3: Decompositions by $N_{\text{ex}}$} (left) and by $\mathrm{SU}(3)$ jointly with $N_{\text{ex}}$} (right), for the normal (bottom) and intruder (top) $3/2^{-}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, for the Daejeon16 interaction, in calculations with $N_{\mathrm{max}}=8$ and ${\hbar\omega}=12.5\,{\mathrm{MeV}}$. The $\mathrm{SU}(3)$ decompositions are shown arranged by the Bohr deformation variables corresponding to the given $\mathrm{SU}(3)$ quantum numbers, and include contributions from the $0{\hbar\omega}$ (blue, shaded circles) and $2{\hbar\omega}$ (red, open circles) spaces. Contributions from irreps which are degenerate with respect to the $\mathrm{SU}(3)$ Casimir operator (connected by dotted lines) cannot be distinguished (and, for plotting purposes, such contributions have, arbitrarily, been distributed equally between these irreps).
Case study: $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}.
The nucleus 11Li provided an early and dramatic exemplar of halo nuclei [53]. Neutron knockout experiments show nearly equal contributions of neutron $1s_{1/2}$ and $0p_{1/2}$ orbitals [54], providing strong evidence for a ground state comprised of a nearly equal mix of normal and intruder states. 11Li is well known to be a challenge to describe fully microscopically, i.e., in terms of three protons and eight neutrons, for example in the NCSM [55, 56, 57], or using fermionic molecular dynamics (FMD) [58]. The FMD calculation addresses mixing of $(p_{1/2})^{2}$ and $(s_{1/2})^{2}$ configurations, that is, normal and intruder states. The nucleus 11Li has also been approximated as a three-body system, typically $^{9}\mathrm{Li}+2n$ [59, 60, 61, 62, 63, 64, 65, 66, 67, 68], allowing $(p_{1/2})^{2}$-$(s_{1/2})^{2}$ mixing to be investigated.
Our calculated excitation spectrum is shown in Fig. 2, as a function of $N_{\mathrm{max}}$, for both the chiral N3LO and Daejeon16 interactions (at fixed basis parameter ${\hbar\omega}=18\,{\mathrm{MeV}}$ and $12.5\,{\mathrm{MeV}}$, respectively). This figure highlights the different evolution with $N_{\text{max}}$} for normal (blue, shaded symbols) and intruder (red, open symbols) states. An illustration of how we may identify normal and intruder states from their NCSM wave functions [35, 37, 39] is provided in Fig. 3 (left), where we display the decomposition of the wave functions into ${N_{\text{ex}}}{\hbar\omega}$ components, in particular, for the $3/2^{-}$ levels.
The $0{\hbar\omega}$ shell model picture for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} gives just $3/2^{-}$ and $1/2^{-}$ states, corresponding to finding the valence proton in the $p_{3/2}$ and $p_{1/2}$ orbitals, respectively. Indeed, the two lowest calculated states (blue, shaded symbols in Fig. 2) match this expectation.
From their decomposition into configurations with different $N_{\text{ex}}$}, these states are clearly normal, as illustrated for the $3/2^{-}$ ground state in Fig. 3 (bottom left). Namely, the single largest contribution to the probability comes from ${N_{\text{ex}}}=0$ (in this case, a single configuration, the filled neutron $0p$ shell), although, naturally for a no-core calculation (see, e.g., Fig. 8 of Ref. [37]), this $0{\hbar\omega}$ configuration is “dressed” with a tail of smaller contributions from excited configurations.
Then, several intruder states arise at higher energy in the calculation (red, open symbols in Fig. 2), starting with a closely spaced $3/2^{-}$ and $1/2^{-}$. (There are many more such states, of which we only show the first $7/2^{-}$ and $5/2^{-}$ as representative examples.) While these start at high energy, for low $N_{\text{max}}$}, they rapidly descend with increasing $N_{\text{max}}$}.
The intruder nature of the excited $3/2^{-}$ and $1/2^{-}$ states is illustrated for the excited $3/2^{-}$ state in Fig. 2 (top left). Note the negligible $0{\hbar\omega}$ contribution, of a few percent, while the strongest contribution comes from $2{\hbar\omega}$ configurations, which are again dressed with a tail of more highly excited oscillator configurations.
Because Daejeon16 is a “softer” interaction, it is perhaps unsurprising that the intruder states found in the Daejeon16 calculation [Fig. 2(b)] descend in energy much sooner (with increasing $N_{\mathrm{max}}$) than those found in the N3LO calculation [Fig. 2(a)]. They also start to level off in energy, suggesting they are approaching the values they would attain in an untruncated calculation.
As an intruder state approaches the normal state of the same angular momentum and parity, these states can mix. In the Daejeon16 calculations [Fig. 2(b)], it is the $1/2^{-}$ normal and intruder states (upward triangles) which approach in energy first, due to the high excitation energy of the normal $1/2^{-}$ state (and relatively smaller spacing between the intruder $3/2^{-}$ and $1/2^{-}$ states). The excited $1/2^{-}$ state initially tracks the other intruder states in energy, but then appears to hit a floor at higher $N_{\text{max}}$}, while the other intruder states continue to fall in energy. On the other hand, the normal $1/2^{-}$ state is pushed down in excitation energy for higher $N_{\text{max}}$}. The two levels approach to within $\approx 2\,{\mathrm{MeV}}$, but no closer. This behavior is suggestive of the level repulsion arising in an avoided crossing [69, 70]. Such a mixing picture is borne out, e.g., by examination of $N_{\text{ex}}$} decompositions (similar to those in Fig. 3, but not shown), which suggest that the normal and intruder states $1/2^{-}$ are nearly maximally mixed by ${N_{\text{max}}}=12$. Although the normal and intruder $3/2^{-}$ states (squares) do not yet approach as closely at the $1/2^{-}$ states in these calculations, we can anticipate that they will likewise strongly mix, as the intruder state $3/2^{-}$ continues to approach the normal $3/2^{-}$ “ground state”.
Figure 4: Dimensionless ratio $Q/r^{2}$ for normal (blue, shaded symbols) and intruder (red, open symbols) $3/2^{-}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}, for the protons (left) and neutrons (right), calculated with the chiral N3LO (top) and Daejeon16 (bottom) interactions. Calculated values are shown as functions of $N_{\text{max}}$}, at fixed $\hbar\omega$} (as indicated). The experimental value for $Q_{p}/r_{p}^{2}$ [71, 72] for the $3/2^{-}$ ground state is shown for comparison (solid square), with the point-proton radius deduced from the charge radius as detailed in Ref. [73].
Now we turn to the role of deformation in the intruder states, and thus, ostensibly, after mixing, in the $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} halo ground state. The electric quadrupole moment is strongly sensitive to deformation. Here we take the dimensionless ratio $Q/r^{2}$, i.e., the quadrupole moment normalized to the mean square radius (or monopole moment). This ratio is more rapidly convergent in NCSM calculations than either the quadrupole moment or mean square radius individually [57, 73] and is the more appropriate measure of deformation for an axially symmetric rotational nucleus [74, 73]. Electromagnetic measurements of the quadrupole moment (and charge radius) are, of course, are only directly sensitive to the proton density distribution. However, in Fig. 4, we examine this ratio both for the proton observables (and thus proton deformation) and for the neutron observables (and thus neutron deformation).
Take first the proton ratio $Q_{p}/r_{p}^{2}$ [Fig. 4 (left)]. (Note that the calculated quadrupole moment is negative, for both the ground and excited $3/2^{-}$ states, so increasing magnitude goes downward in the plot.) For the ground state, a negative sign for the quadrupole moment is as expected for a single nucleon outside a closed shell (e.g., Refs. [75, 76]). For the excited $3/2^{-}$ state, if it is taken to be a member of a $K=1/2$ band (with inverted spin orderings due to Coriolis staggering), such a negative (spectroscopic) quadrupole moment might be associated with prolate intrinsic deformation (e.g., Fig. 1 of Ref. [77]), at least, to the extent that it might be meaningful to associate a “deformation” to the density distribution generated by just three protons. Regardless, the proton $Q/r^{2}$ is nearly the same for the normal and intruder states ($Q_{p}/r_{p}^{2}\approx-0.6$); this at least suggests the proton structure is not very different between the two. Our results, in the calculations with either the N3LO or Daejeon16 interaction, are in close agreement with experiment [71, 72] (solid square in Fig. 4). (Previous NCSM calculations obtained similar values of $Q_{p}/r_{p}^{2}$ [55, 56], including already with the Daejeon16 interaction [57].)
The calculated neutron quadrupole moment ratio $Q_{n}/r_{n}^{2}$ [Fig. 4 (right)] is much smaller for the normal state ($Q_{n}/r_{n}^{2}\approx-0.2$), as one might expect for small deviations from a spherical filled neutron shell, but much larger for the intruder state ($Q_{n}/r_{n}^{2}\approx-1.3$), as expected for a more highly deformed state arising from excitations of neutrons across the shell closure. At least, this is the behavior of the observables for the N3LO calculations [Fig. 4 (b)].
However, in the Daejeon16 calculations [Fig. 4 (d)], the situation is more subtle, as we might expect in the presence of mixing. For the ground state, the neutron quadrupole moment starts small, at low $N_{\text{max}}$}, as in the N3LO calculation, but then it steadily climbs, from ${N_{\text{max}}}=4$ onward. For the excited $3/2^{-}$ state, the neutron quadrupole moment attains a large value at low $N_{\text{max}}$}, as in the N3LO calculation, but then turns over and starts declining in magnitude after ${N_{\text{max}}}=8$. This reflects an intruder component (with large neutron deformation) mixing into the ground state, and a normal component (with negligible neutron deformation) mixing into the excited state.
To proceed beyond the simple measure of deformation provided by $Q/r^{2}$ and more properly characterize the “shape” of the state, we decompose the wave function into components in irreps of Elliott’s $\mathrm{SU}(3)$ [51] and interpret them approximately in terms of the standard quadrupole shape parameters $\beta$ and $\gamma$ [52]. In Fig. 3 (right), we display the structure of the normal and intruder $3/2^{-}$ wave functions, decomposed first in terms of excitation quanta $N_{\mathrm{ex}}$, and then, within each ${N_{\text{ex}}}\hbar\omega$ space, by $\mathrm{SU}(3)$ irreps $(\lambda,\mu)$. These are arranged by their location on the $(\beta,\gamma)$ plane, where the area of each circle is proportional to the probability (i.e., contribution to the wave function norm) from that given $N_{\text{ex}}$} and $(\lambda,\mu)$. (Specifically, these results are for the $N_{\mathrm{max}}=8$ calculation using Daejeon16 at $\hbar\omega=12.5$ MeV; the plots for higher values of $N_{\mathrm{max}}$ are not very different.) As already noted [Fig. 3 (left)], the normal $3/2^{-}_{1}$ state (in this calculation) is largely $0{\hbar\omega}$, while the $0{\hbar\omega}$ contribution to the intruder $3/2^{-}_{2}$ is less than $5\%$.
Moreover, the shapes of these states differ significantly. The normal $3/2^{-}_{1}$ state is dominated by the nearly spherical unique $0{\hbar\omega}$ irrep $(1,0)$ of $\mathrm{SU}(3)$ (blue, shaded circle in the $\beta$-$\gamma$ plot), while the intruder $3/2_{2}^{-}$ state has its largest contributions from highly deformed $2{\hbar\omega}$ $(5,2)$ and $(6,0)$ irreps (open circles in $\beta$-$\gamma$ plot). The $(6,0)$ irrep represents a prolate, axially symmetric shape (or as close to this as can be attained in the $2{\hbar\omega}$ model space for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}), but the $(5,2)$ irrep lies just to the prolate side of the dotted line indicating maximal triaxiality ($\gamma=30^{\circ}$). Thus, the $\mathrm{SU}(3)$ decomposition in Fig. 3 (top right) indicates significant deviations from axial symmetry for the intruder state.
Thus, despite the numerically large calculation and a priori complex nature of the microscopic wave functions, the final picture is fairly simple: the normal state follows the naive spherical shell-model expectations, while the intruder is both mostly $2{\hbar\omega}$ (ostensibly two-particle, two-hole) and well-deformed.
Figure 5: Relative energies for the normal $1/2+$ and $5/2+$ (blue, shaded symbols) and first intruder $1/2+$ through $5/2+$ (red, open symbols) levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, for (a) a chiral N3LO interaction and (b) the Daejeon16 interaction, shown as functions of $N_{\mathrm{max}}$ (at fixed $\hbar\omega$}, as indicated).
Refer to captionRefer to caption
Figure 6: Decompositions by $N_{\text{ex}}$} (left) and by $\mathrm{SU}(3)$ jointly with $N_{\text{ex}}$} (right), for the normal (bottom) and intruder (top) $5/2^{+}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, for the Daejeon16 interaction, in calculations with $N_{\mathrm{max}}=6$ and ${\hbar\omega}=17.5\,{\mathrm{MeV}}$. The $\mathrm{SU}(3)$ decompositions shown arranged by the Bohr deformation variables corresponding to the given $\mathrm{SU}(3)$ quantum numbers, and include contributions from the $0{\hbar\omega}$ (blue, shaded circles) and $2{\hbar\omega}$ (red, open circles) spaces. Contributions from irreps which are degenerate with respect to the $\mathrm{SU}(3)$ Casimir operator (connected by dotted lines) cannot be distinguished (and, for plotting purposes, such contributions have, arbitrarily, been distributed equally between these irreps).
Case study: $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}.
The nuclide 29F may be viewed as a cousin to 11Li, shifted up by a major oscillator shell. It is another halo nucleus [78] on an island of inversion [79]. Based at least on model arguments, though not direct experimental evidence, it is also thought to be deformed [80, 81]. Our maximum model space is only $N_{\mathrm{max}}=6$, with an $M$-scheme dimension of $4.7\times 10^{9}$. Nonetheless we can see trends which closely parallel those found above for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$}.
The calculated energy levels, for the same two interactions (chiral N3LO and Daejeon16) are shown in Fig. 5, again highlighting their evolution with $N_{\text{max}}$}, at fixed $\hbar\omega$} (here, ${\hbar\omega}=18\,{\mathrm{MeV}}$ and $17.5\,{\mathrm{MeV}}$, respectively). The normal” levels (blue, shaded symbols) come lowest, as expected. These are now a $5/2^{+}$ ground state and $1/2^{+}$ excited state, corresponding to finding the valence neutron in the $0d_{5/2}$ and $1s_{1/2}$ orbitals, respectively (the $3/2^{+}$ level corresponding to the $0d_{3/2}$ orbital lies somewhat higher, at $\gtrsim 10\,{\mathrm{MeV}}$, and has been omitted for clarity). Then several more closely spaced intruder levels (red, open symbols) start at much higher energy, but swoop down with increasing $N_{\text{max}}$}. The content of the wave functions in terms of oscillator excitations, shown in Fig. 6 (left) for the $5/2^{+}$ levels, resembles that seen above for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} in Fig. 3 (left) — a dominant contribution at $0{\hbar\omega}$ or $2{\hbar\omega}$ respectively, for normal and intruder states, but dressed with a decaying tail of contributions at higher $N_{\text{ex}}$}.
However, once again, the intruder levels begin not quite so high in energy when calculated with the softer Daejeon16 interaction [Fig. 5(b)] — in fact, the starting energy of $\approx 13\,{\mathrm{MeV}}$, for ${N_{\text{max}}}=2$, is far below the energy of twice $\hbar\omega$} one might naively expect from the shell model. And, once again, the intruder levels approach the low-lying normal levels much sooner with the Daejon16 interaction. The intruder $1/2^{+}$ is the first to closely approach its normal counterpart (upward triangles). Indeed, already by ${N_{\text{max}}}=6$, the two approach to within $\approx 2\,{\mathrm{MeV}}$. Such strong mixing as was found for the $1/2^{-}$ states in $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} at similar separation [Fig. 2(b)] is not yet seen here, suggesting a smaller mixing matrix element ($1\,{\mathrm{MeV}}$). Nonetheless, as the intruder energies continue to fall with $N_{\text{max}}$}, we can anticipate mixing to occur, or perhaps even an outright crossing yielding an intruder ground state.
The quadrupole moments for the normal and intruder $5/2^{+}$ levels, taken again in the dimensionless ratio $Q/r^{2}$, as a measure of deformation, are shown in Fig. 7. The pattern here is, again, markedly similar to that seen above for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} in Fig. 4. For the proton quadrupole moments (left), there is very little difference between the normal (blue, shaded symbols) and intruder (red, open symbols) ratios, suggesting the proton structure is nearly the same (this ratio is, again, not necessarily to be interpreted as yielding a deformation, given just a single proton outside a closed shell). Then, for the neutron quadrupole moment (right), the normal neutron quadrupole moment is again nearly zero, as expected for a filled spherical shell, and large for the intruder state, suggesting strong deformation.
The picture of a spherical normal state and a deformed intruder is supported by the decomposition into $\mathrm{SU}(3)$ irreps, shown in Fig. 6 (right), again interpreted in terms of Bohr deformation parameters. The normal state [Fig. 6 (bottom right)] is, necessarily, dominated by the unique $0{\hbar\omega}$ irrep $(2,0)$, corresponding to a single nucleon in the $sd$ shell. The decomposition of the intruder state [Fig. 6 (top right)] is more fragmented than for $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} [Fig. 3 (top right)]. The strongest contributions are from the highly deformed $2{\hbar\omega}$ $(9,2)$ and $(10,0)$ irreps, but the probablility is spread over several of the most highly deformed $2{\hbar\omega}$ irreps. These again cover a range of triaxiality, all on the prolate side of $\gamma=30^{\circ}$. The spread over irreps is perhaps unsurprising, considering the lore that spin-orbit splitting, which breaks $\mathrm{SU}(3)$, increases with mass number $A$. Moreover, from comparing Fig. 6 with Fig. 3, it is apparent that there is simply a greater density of possible $\mathrm{SU}(3)$ irreps for the nucleus to “choose” between, within a similar region of the Bohr $\beta$-$\gamma$ deformation space.
While taking $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$} to higher $N_{\text{max}}$}, closer to convergence, is beyond current capabilities, we already have strong evidence that the lessons learned from $^{\makebox[7.97226pt][r]{$\scriptstyle 11$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{Li}$} apply here as well. We confirm both the importance and the numerical challenge of intruders for these island-of-inversion nuclides.
Figure 7: Dimensionless ratio $Q/r^{2}$ for the normal (blue, shaded symbols) and intruder (red, open symbols) $5/2^{+}$ levels of $^{\makebox[7.97226pt][r]{$\scriptstyle 29$}}_{\makebox[7.97226pt][r]{$\scriptstyle$}}\mathrm{F}$}, for the protons (left) and neutrons (right), calculated with the Daejeon16 interaction. Calculated values are shown as functions of $N_{\text{max}}$} (at fixed $\hbar\omega$}, as indicated).
Summary.
We have carried out large no-core shell model calculations of the island-of-inversion nuclides 11Li and 29F, using two interactions, a chiral N3LO interation and the phase-shift-equivalent interaction Daejeon16. In a naive spherical shell model picture, these nuclides should have very simple structure: spherical filled neutron shells and one proton outside filled proton shells. Instead, the ground state and other low-lying states are inverted to be primarily, or at least strongly mixed with, deformed two-particle, two-hole intruders. In ab initio NCSM calculations, while the intruders descend towards the spherical normal states with increasing model space ($N_{\text{max}}$}), convergence is slow, even for the softer Daejeon16 interaction, and thus the intruders are a challenge to reproduce in the NCSM. Interestingly, the intruder states may be “simple” in a group-theoretical framework, that is, dominated by a few irreps, representing the largest possible deformations with some triaxial tendencies).
Confronted with the paradox of the simple naive spherical picture versus the more complicated deformed reality, at least for the spherical shell model, we propose these nuclides as important test cases for advanced nuclear structure methods. Examples include the symmetry-adapted configuration-interaction calculations [82, 83, 84, 85], the recently introduced angular-momentum-projected coupled cluster method [86, 87], or perhaps novel methods, yet to be proposed, in which these “intruder” configurations become the new normal.
Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Numbers DE-FG02-95ER40934 and DE-FG02-03ER41272. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231, using NERSC awards NP-ERCAP0023497 and NP-ERCAP0031993. Part of this research was enabled by computational resources supported by a generous gift to SDSU from John Oldham.